A weather balloon rises 150 meters the first minute after it is released. Each minute, it rises 6% slower than the previous minute, How far does the balloon travel in 15 minutes? Give your answer to the nearest tenth.
Answer by amoresroy(361)

a = 150 meters
r = .94 (6% lower)
n = 15
S = travel distance of balloon in 15 mins
S = a (1-r^n)/ (1-r)
= 150 (1-.94^15) / (1-.94)
= 150 (10.07847
= 1511.8
The balloon travels 1511.8 meters in 15 mins.